D Question 12 1 pts An IV piggyback containing 100 mL NS is to infuse over 30 minutes q.6h. How many mL/hr will the nurse set on the IV infusion pump? D Question 13 3 pts The MD has ordered the following IVs to run consecutively over 24 hrs. #1 #2 #3 925 mL RL 750 mL DNS 600 mL NS c 20 mEq KCl 1. Calculate the volume per hour 2. Calculate the total volume per minute using standard macrodrip tubing 3. Calculate the total volume per minute using standard microdrip tubing
Answer #1

12 . The volume of NS to be infused = 100 ml

Total time to infuse = 30 mins

Flow rate = total volume ( ml) / total time ( hr)

60 mins = 1hr

Then 30 mins = 1/2 hr

Flow rate = 100 × 1 / 2

= 100 / 2

= 50 ml / hr .

13. 1) # Ordered dose of RL is 925 ml, to be infused in 24hr

Flow rate will be = total volume / total time

= 925ml / 24 hr

= 38.5 ml / hr is volume per hour

# Ordered dose of D5NS = 750 ml over 24 hr

Volume per hour = total volume / total time

= 750 ml/ 24 hr

= 31.25 ml / hr

# Ordered dose is 600ml NS with 20 mEq kcl , to be infused in 24hour

Volume per hour = total volume / total time

= 600ml/ 24hr

= 25ml / hr

2) # Volume per minute using standard macrodrip tubing is

Total volume of RL = 925 ml

Total time is 24hours = 24× 60 mins

Drop factor for macrodrip tubing is 20

Volume / min = 925ml/ 24×60 × 20

= 12.8 ml / min

# 750 ml of D5NS to be I fused in 1440 mins

Drop factor of macrodrip set = 20

Volume / min = 750 ml / 1440 × 20

= 10.41 ml/ mins

# 600 ml of NS with 20mEq kcl over 1440 mins, drop factor 20 using macrodrip set

Volume / min = 600 / 1440 × 20

= 8.33 ml / min

3) # Volume / minute using a standard microdrip tubing

Total volume of RL = 925 ml

Total time in min = 1440 ( 24× 60)

Drop factor of microdrip set = 60

Volume / min = 925 / 1440 × 60

= 38.54ml / min

# Volume of D5NS is 750 ml

Total time to infuse = 1440min

Drop factor = 60

Volume / min = 750 / 1440 × 60

= 31.25 ml / min

#Volume of NS to be infused = 600 ml

Total time to infuse = 1440min

Drop factor of microdrip set = 60

Volume per min = 600/1440 × 60

= 25 ml / min